OpenStudy (anonymous):

Find the smallest of three consecutive integers such that 8 more than the product of the first and second is equal to the product of the second and third.

OpenStudy (anonymous):

8+ab=bc a+1=b b+1=c solve the system of equations

OpenStudy (agreene):

8+(n*(n+1))=(n+1)(n+2) solve for n where n is an integer and your three consecutive integers are n, n+1 and n+2.